OpenStudy (anonymous):

Find all solutions in interval [0,2pi) Sec theta + sq.root 2 = 0 Find solution (s) in terms of pi (radians)

OpenStudy (tkhunny):

Have you considered \(\sec(\theta) = 1/\cos(\theta)\)?

OpenStudy (anonymous):

for answers i got 3pi/4 and 5pi/4 but im not sure thats right

OpenStudy (tkhunny):

How did you get that? \(\cos(\theta) = -\dfrac{\sqrt{2}}{2}\) Look sto me like you have it. Why do you doubt?

OpenStudy (anonymous):

\[\theta =\sec ^{-1}\left(-\sqrt{2}\right)=\frac{3 \pi }{4} \]